Determine the concentration of a solution (in % by mass, % by volume, % by mass

volume, molarity, molality and ppm) containing 5 grams of NaCI and 50 ml of water.(p of NaCl = 2.1 g/ml,p of water = 1g/ml).What is the mole fraction of NaCI in the solution?


1
Expert's answer
2021-05-20T07:53:54-0400

Ans:-

Given "5" grams of "NaCI" and "50 ml" of water

Moles of "NaCl=\dfrac{5}{58.4}=0.0856mol"


Concentration of Solution in molarity"=\dfrac{Moles of Sollution}{volume of solution}=\dfrac{0.0856}{50}\times1000=1.7123M"


Concentration of Solution in molality"=\dfrac{moles of solute}{Mass of Solvent}=\dfrac{0.0856}{50}\times1000=1.7123m"


We know that density of Water is "1g\/ml".

So, mass of Water present in solution "=" density"\times" volume"=1\times50=50g"

Hence moles of water"=\dfrac{50}{18}=2.77mol"

Mole fraction of "NaCl" in solution"= \dfrac{0.0856}{0.0856+2.77}=0.03"